OpenStudy (anonymous):

16 people get an order of 20 breadsticks. How many breadsticks should each person get?

OpenStudy (amistre64):

(Total amount)/(number of people) (20)/(16) = 10/8 = 5/3 = 1 stick plus (2/3) of another stick per person

OpenStudy (anonymous):

correction: 10/8 = 5/4 so 1 stick plus (1/4) of another stick per person